# 2Departament de Física Quàntica i Astrofísica and Institut de Ciències del Cosmos, Universitat de Barcelona, 08028 Barcelona, Spain

### v1

@@ Abstract
We investigate the scattering of electromagnetic and gravitational waves off a Reissner-Nordström black hole in the low-temperature regime where the near-horizon throat experiences large quantum fluctuations. We find that the black hole is transparent to electromagnetic and gravitational radiation below a certain frequency threshold. This phenomenon arises because the angular momentum of the black hole is quantized, creating an energy gap between the spinless black hole state and the first excited spinning states. Radiation with angular momentum—such as photons, gravitons, and partial waves of a massless scalar field, which we also study—must supply enough energy to bridge this gap to be absorbed. Below this threshold, no absorption can occur, rendering the black hole transparent. For frequencies above the gap, the scarcity of black hole states continues to suppress the absorption cross-section relative to semiclassical predictions, making the black hole translucent rather than completely transparent. Notably, electromagnetic absorption is significantly stronger than gravitational absorption, beyond what differences in spin alone would suggest.
@@ 1 Introduction and Summary
In this article, we study the absorption of electromagnetic and gravitational waves—the two massless radiation fields known to exist in nature—by a near-extremal Reissner-Nordström black hole. We find that, at sufficiently low frequencies, the absorption cross-section vanishes and the black hole becomes transparent. This is in sharp contrast with the classical view of black holes as perfect absorbers: quantum effects render them invisible to light and gravitational waves below a certain frequency threshold.
@@
For frequencies above the bound (1.3), the absorption of spinning waves becomes nonzero but remains small, as the density of accessible black hole states is still low. This happens despite the fact that, in the quantum regime, transition rates between individual states are enhanced, reflecting the general feature that late-time correlations in a quantum black hole decay as a power law, rather than exponentially as in the semiclassical case. However, unlike the s-wave case studied in Emparan:2025sao , here this enhancement is outweighed by the scarcity of absorbing black hole states across the entire frequency range where comparison with semiclassical results is meaningful. The black hole, though not entirely transparent to spinning radiation, remains very faint—i.e., translucent—and does not recover its opacity until the absorbed energy becomes large enough to push it out of the quantum regime. Among other consequences, transparency and translucency should affect the phenomenon of black-hole-induced decoherence Danielson:2022tdw ; Biggs:2024dgp ; Li:2025vcm .
@@
In this article we focus on charged black holes in four-dimensional Einstein-Maxwell theory, but the results extend naturally to higher-dimensional charged black holes. A similar transparency to photons and gravitons is also expected in near-extremal neutral rotating black holes, such as the Kerr solution.
@@ 2 Setup
We follow the framework used in Emparan:2025sao for the absorption of waves by a four-dimensional Reissner-Nordström black hole, now extended to spinning waves with $\ell\geq 1$. We will borrow results about the interaction of the black hole with various fields from Brown:2024ajk , which studied it with a view to spontaneous Hawking emission. The technical extension to absorption is simple, but examining the response to incoming radiation places the problem in a different physical context, which brings out new aspects of quantum black hole physics.
@@ 2.1 Scattering waves from infinity into the throat
We consider a monochromatic polarized wave of these fields with frequency $\omega$ and angular momentum $\ell$, which we model as a coherent state of the field with a large occupation number. This wave propagates classically on the background geometry from asymptotic infinity down to the mouth of the black hole throat, where it encounters the black hole as a quantum object. The semiclassical approximation breaks down there, and the propagation of the wave in the throat geometry is replaced by the interaction between the wave and the quantum states of the black hole described by the Schwarzian theory. This interaction accounts for the quantum backreaction of the black hole to the incoming radiation.
@@
[…]
@@ 2.2 Transition rates
When the perturbation is weak we can apply Fermi’s Golden Rule, which gives us the absorption and stimulated emission transition rates in the form
@@
where $|i\rangle$ and $|f\rangle$ are the initial and final states, with energies $E_{i}$ and $E_{f}$ (which are always measured above extremality), and where $\rho(E_{f})$ is the density of final states. [^6] To apply Fermi’s Golden Rule, the interaction between the incoming wave and the quantum Schwarzian must last a time $t$ such that $1/(E_{i}\pm\omega)\ll t\ll 1/|\langle f|V_{0}\,\mathcal{O}_{\Delta}|i\rangle|$, where the plus sign is for absorption and the minus sign for stimulated emission. The first bound guarantees that only the final states with $E_{f}\approx E_{i}\pm\omega$ contribute to the transition probability. The second bound ensures that the transition probability always remains small (i.e., that our perturbation remains weak). For the parameter range we are interested in, such a time window always exists, and it is typically extremely large, thanks to the exponential $e^{-S_{0}}$ in the matrix elements (3.3). As in Emparan:2025sao , we can treat the radiation states as eigenstates of the occupation number instead of coherent states, since for large occupation numbers their variance is negligible. For the same reason, we can neglect spontaneous emission, i.e., Hawking radiation.
@@
To compute the total absorption cross-section for a given partial wave of angular momentum $\ell$ with fixed polarization, we must average over the initial angular momentum states $m_{\text{\tiny BH}}$, and sum over the possible values of the final angular momentum $j^{\prime}$ of the black hole, as well as over the quantum numbers $m^{\prime}_{\text{\tiny BH}}$ and $m$.
@@ 3 Computing the cross-section
bearing in mind that our initial spinless black hole has $m_{\text{\tiny BH}}=0$.
@@
Here we use again the convention that denotes a product of Pochhammer symbols for the four choices of sign. These expressions are convenient since the conformal dimension $\Delta$ is entirely contained in the factors ${B}^{\Delta}_{E_{i}\pm\omega}$ that multiply the term (3.10), which gave the transition rates in Emparan:2025sao for $\Delta=1$.
Now, for $E_{f}=E_{i}\pm\omega$ the expression (3.6) of $\mathcal{A}^{\Delta}_{E_{f}}$ reduces to
@@
We can now derive the semiclassical limit $\omega,E_{b}\ll E_{i}$ of the quantum greybody factor (3.7). We find
$P_{\text{gb}}\approx|\mathcal{N}_{\ell}|^{2}\,\,\frac{2^{2\Delta+1}\,\pi^{2\Delta}\,(2\ell+1)}{\Gamma(2\Delta)}\,\frac{r_{0}^{2(\Delta+\ell)}\,\omega^{2\ell+1}}{\beta^{2\Delta-1}}\,\sinh\!\left(\!\frac{\beta\omega}{2}\!\right)\,\Gamma\!\left(\!\Delta\pm i\,\frac{\beta\omega}{2\pi}\!\right)\,,$ (3.17)
where we have used the identity
$\sinh\!\left(\!\frac{\beta\omega}{2}\!\right)\,\Gamma\!\left(\!\Delta\pm i\,\frac{\beta\omega}{2\pi}\!\right)=\frac{2\pi^{2}}{\beta\omega}\,\left(i\,\frac{\beta\omega}{2\pi}\right)_{\Delta}\,\left(-i\,\frac{\beta\omega}{2\pi}\right)_{\Delta}$ (3.18)
to express the Pochhammer symbols in (3.7) in terms of Gamma functions. The inverse temperature near extremality is given by
@@
It is also illustrative to take the very low-frequency limit, $\omega\ll E_{b}\ll E_{i}$, where $\omega\ll\beta^{-1}$. Our result (3.17) simplifies to
@@ 4.1 Quantum spin transparency
In this case there is no obstacle to absorption for any frequency $\omega>0$, while stimulated emission will occur only if it takes the black hole down to a state above the spin gap, i.e., $\omega<E_{i}-E_{0,\ell}$. When no spin is involved, i.e., $\ell=0$, stimulated emission is restricted to $\omega<E_{i}$, which leads to the kinks in Figure 1 of Emparan:2025sao for the absorption cross-section as a function of $\omega$ [^7] We take this opportunity to correct a typo in Figs. 2 and 3 of Emparan:2025sao : the caption states $E_{i}=0.1\,E_{b}$, but the plots actually correspond to $E_{i}=0.01\,E_{b}$.
@@
In the analysis and figures below, we focus on this regime with $\ell\geq 1$ and $E_{i}<E_{0,\ell}$, where the distinctive quantum features of transparency and absence of stimulated emission emerge. The spin-induced gap also closes the emission channel for single quanta, suppressing spontaneous emission. This led Brown:2024ajk to consider spin-singlet pairs of entangled photons or gravitons, which in our setup are not stimulated.
@@ 4.2 Results for $\ell=1$ photon and $\ell=2$ graviton
[…]
$-\frac{\left(4E_{b}^{2}+\omega^{2}+4E_{b}(2E_{i}-\omega)\right)\left(E_{b}^{2}+4\omega^{2}+4E_{b}(2E_{i}-\omega)\right)\sinh\!\left(\!2\pi\sqrt{\frac{2(E_{i}-\omega-E_{b})}{E_{b}}}\right)\Theta(E_{i}-\omega-E_{b})}{\cosh\!\left(\!2\pi\sqrt{\frac{2E_{i}}{E_{b}}}\right)-\cosh\!\left(\!2\pi\sqrt{\frac{2(E_{i}-\omega)}{E_{b}}}\right)}\Bigg{]}$ (4.6)
@@
[…]
$-\frac{\left(E_{b}^{2}+4\omega^{2}+4E_{b}(2E_{i}-\omega)\right)\sinh\!\left(\!2\pi\sqrt{\frac{2(E_{i}-\omega-3E_{b})}{E_{b}}}\right)\Theta(E_{i}-\omega-3E_{b})}{\cosh\!\left(\!2\pi\sqrt{\frac{2E_{i}}{E_{b}}}\right)-\cosh\!\left(\!2\pi\sqrt{\frac{2(E_{i}-\omega)}{E_{b}}}\right)}\Bigg{]}\,.$ (4.7)
@@
Figure 3: Field-type-independent absorption cross-section $\widehat{\sigma}$ (cf. (4.8)) as a function of $\omega/E_{i}$ for $E_{i}=0.8\,E_{b}$. Left: plot at fixed $\ell=1$ for different $\Delta$. Right: plot at fixed $\Delta=3$ for different $\ell$. In the quantum regime where $\omega/E_{i}\lesssim\mathcal{O}(1)$ (say, $\omega/E_{i}\leq 5$), but above the transparency window for the $\ell=2$ graviton, which ends at $\omega/E_{i}=2.75$, the cross-section becomes higher as we increase $\Delta$, and as we decrease $\ell$. The same qualitative feature is observed for fixed $\ell=2$, or for fixed $\Delta=2$. The length unit for $\widehat{\sigma}$ is $E_{b}^{\Delta+\ell-1}\,r_{0}^{\Delta+\ell}$.

### v3

# 2Departament de Física Quàntica i Astrofísica and Institut de Ciències del Cosmos, Martì i Franquès 1, Universitat de Barcelona, 08028 Barcelona, Spain

@@ Abstract
We investigate the scattering of electromagnetic and gravitational waves off a Reissner-Nordström black hole in the low-temperature regime where the near-horizon throat experiences large quantum fluctuations. We find that the black hole is transparent to electromagnetic and gravitational radiation of fixed helicity below a certain frequency threshold. This phenomenon arises because the angular momentum of the black hole is quantized, creating an energy gap between the spinless black hole state and the first excited spinning states. Radiation with angular momentum—such as photons, gravitons, and partial waves of a massless scalar field, which we also study—must supply enough energy to bridge this gap to be absorbed. Below this threshold, no absorption can occur, rendering the black hole transparent. For frequencies above the gap, the scarcity of black hole states continues to suppress the absorption cross-section relative to semiclassical predictions, making the black hole translucent rather than completely transparent. Notably, electromagnetic absorption is significantly stronger than gravitational absorption, beyond what differences in spin alone would suggest.
@@ 1 Introduction and Summary
In this article, we study the absorption of electromagnetic and gravitational waves—the two massless radiation fields known to exist in nature—by a near-extremal Reissner-Nordström black hole. We find that, at sufficiently low frequencies, the absorption cross-section for a coherent, circularly polarized partial wave vanishes and the black hole becomes transparent. [^2] If the wave is unpolarized, or linearly polarized, it can be absorbed in the form of spin-singlet di-quanta. However, this absorption is a strongly suppressed higher-order phenomenon Brown:2024ajk . This is in sharp contrast with the classical view of black holes as perfect absorbers: quantum effects render them invisible to light and gravitational waves below a certain frequency threshold.
@@
For frequencies above the bound (1.3), the absorption of spinning waves becomes nonzero but remains small, as the density of accessible black hole states is still low. This happens despite the fact that, in the quantum regime, transition rates between individual states are enhanced, reflecting the general feature that late-time correlations in a quantum black hole decay as a power law, rather than exponentially as in the semiclassical case. However, unlike the s-wave case studied in Emparan:2025sao , here this enhancement is outweighed by the scarcity of absorbing black hole states across the entire frequency range. The black hole, though not entirely transparent to spinning radiation, remains very faint—i.e., translucent—and does not recover its opacity until the absorbed energy becomes large enough to push it out of the quantum regime. Among other consequences, transparency and translucency should affect the phenomenon of black-hole-induced decoherence Danielson:2022tdw ; Biggs:2024dgp ; Li:2025vcm .
@@
In this article we focus on charged black holes in four-dimensional Einstein-Maxwell theory, but the results extend naturally to higher-dimensional charged black holes. A similar transparency to photons and gravitons is also expected in near-extremal neutral rotating black holes, such as the Kerr solution. [^5] Unfortunately, the prospects of observing this effect in astrophysical black holes are vanishingly small. The scale of the relevant wavelengths is $\sim r_{0}S_{0}$, which for a solar-mass black hole is about $10^{80}$ m, or $10^{54}$ times the size of the observable universe. Any such observation would require $\sim 10^{64}$ years, which is impractical.
@@ 2 Setup
We follow the framework used in Emparan:2025sao for the absorption of waves by a four-dimensional near-extremal Reissner-Nordström black hole, now extended to spinning waves with $\ell\geq 1$. We will borrow results about the interaction of the black hole with various fields from Brown:2024ajk , which studied it with a view to spontaneous Hawking emission. The technical extension to absorption is simple, but examining the response to incoming radiation places the problem in a different physical context, which brings out new aspects of quantum black hole physics.
@@ 2.1 Scattering waves from infinity into the throat
We consider a monochromatic, circularly polarized wave of these fields with frequency $\omega$ and angular momentum $\ell$, which we model as a coherent state of the field with a large occupation number. This wave propagates classically on the background geometry from asymptotic infinity down to the mouth of the black hole throat, where it encounters the black hole as a quantum object. The semiclassical approximation breaks down there, and the propagation of the wave in the throat geometry is replaced by the interaction between the wave and the quantum states of the black hole described by the Schwarzian theory. This interaction accounts for the quantum backreaction of the black hole to the incoming radiation.
@@
$\phi_{\text{bdy}}(t)=\mathcal{N}_{\ell}\,\,r_{0}^{\Delta+\ell}\omega^{\ell+1/2}\,\left(A_{\omega\ell m}\,e^{-i\omega t}+\text{h.c.}\right)\,.$ (2.3)
@@ 2.2 Transition rates
When the perturbation is weak we can apply Fermi’s Golden Rule, which gives us the absorption and stimulated emission transition rates in the form [^8] To apply Fermi’s Golden Rule, the interaction between the incoming wave and the quantum Schwarzian must last a time $t$ such that $1/(E_{i}\pm\omega)\ll t\ll 1/|\langle f|V_{0}\,\mathcal{O}_{\Delta}|i\rangle|$, where the plus sign is for absorption and the minus sign for stimulated emission. The first bound guarantees that only the final states with $E_{f}\approx E_{i}\pm\omega$ contribute to the transition probability. The second bound ensures that the transition probability always remains small (i.e., that our perturbation remains weak). For the parameter range we are interested in, such a time window always exists, and it is typically extremely large, thanks to the exponential $e^{-S_{0}}$ in the matrix elements (3.3).
@@
where $|i\rangle$ and $|f\rangle$ are the initial and final states, with energies $E_{i}$ and $E_{f}$ (which are always measured above extremality), $\mathcal{O}_{\Delta}\equiv\mathcal{O}_{\Delta}(0)$, and $\rho(E_{f})$ is the density of final states. [^9] Note that, although the black hole starts in a fixed-energy state, it will evolve through Hawking emission into an ensemble of states in a time $\sim E_{b}^{-1}$, comparable to the duration of our scattering experiment Biggs:2025nzs . As in Emparan:2025sao , we can treat the radiation states as eigenstates of the occupation number instead of coherent states, since for large occupation numbers their variance is negligible. For the same reason, we can neglect spontaneous emission, i.e., Hawking radiation.
@@
To compute the total absorption cross-section for a given partial wave of angular momentum $\ell$ with fixed circular polarization, we must average over the initial angular momentum states $m_{\text{\tiny BH}}$, and sum over the possible values of the final angular momentum $j^{\prime}$ of the black hole, as well as over the quantum numbers $m^{\prime}_{\text{\tiny BH}}$ and $m$.
@@ 3 Computing the cross-section
bearing in mind that our initial spinless black hole has $j=0$, $m_{\text{\tiny BH}}=0$ and $j^{\prime}=\ell$.
@@
Here we use again the convention that denotes a product of Pochhammer symbols for the four choices of sign. The products nicely simplify to a single product of real polynomials in $E_{i}/E_{b}$ and $\omega/E_{b}$,
$\mathcal{B}^{\Delta}_{E_{i}\pm\omega}=\prod_{n=1}^{\Delta-1}\!\left[n^{4}+\frac{4n^{2}\,(2E_{i}\pm\omega)}{E_{b}}+\frac{4\omega^{2}}{E_{b}^{2}}\right]\,.$ (3.15)
These expressions are convenient since the conformal dimension $\Delta$ is entirely contained in the factors ${B}^{\Delta}_{E_{i}\pm\omega}$ that multiply the term (3.10), which gave the transition rates in Emparan:2025sao for $\Delta=1$. Now, for $E_{f}=E_{i}\pm\omega$ the expression (3.6) of $\mathcal{A}^{\Delta}_{E_{f}}$ reduces to
@@
We can now derive the semiclassical limit $\omega,E_{b}\ll E_{i}$ of the quantum greybody factor (3.7). In this regime, the Pochhammers (3.15) reduce to
$\mathcal{B}^{\Delta}_{E_{i}\pm\omega}$ $\approx\prod_{n=1}^{\Delta-1}\!\left(\frac{8n^{2}E_{i}}{E_{b}}+\frac{4\omega^{2}}{E_{b}^{2}}\right)$ (3.18)
$=\left(\!\frac{4\pi}{E_{b}\beta}\!\right)^{2(\Delta-1)}\,\left(\!1+i\,\frac{\beta\omega}{2\pi}\!\right)_{\Delta-1}\,\left(\!1-i\,\frac{\beta\omega}{2\pi}\!\right)_{\Delta-1}$ (3.19)
$=\frac{E_{b}}{2\pi\omega}\,\left(\!\frac{4\pi}{E_{b}\beta}\!\right)^{2\Delta-1}\,\sinh\!\left(\!\frac{\beta\omega}{2}\!\right)\,\Gamma\!\left(\!\Delta+i\,\frac{\beta\omega}{2\pi}\!\right)\,\Gamma\!\left(\!\Delta-i\,\frac{\beta\omega}{2\pi}\!\right)\,$, (3.20)
where the inverse temperature near extremality is given by
@@
In this limit we have
$\Theta(E_{i}\pm\omega-E_{0,\ell})\to 1\,.$ (3.22)
This helps simplify the term $\mathcal{A}^{\Delta}_{E_{i}+\omega}-\mathcal{A}^{\Delta}_{E_{i}-\omega}$ in the cross-section. Up to the Pochhammer factors (3.20), it is now proportional to
$\frac{\sinh\!\left(2\pi\,\sqrt{\frac{2\,(E_{i}+\omega-E_{0,\ell})}{E_{b}}}\right)}{\cosh\!\left(2\pi\,\sqrt{\frac{2(E_{i}+\omega)}{E_{b}}}\right)-\cosh\!\left(2\pi\,\sqrt{\frac{2E_{i}}{E_{b}}}\right)}-\frac{\sinh\!\left(2\pi\,\sqrt{\frac{2\,(E_{i}-\omega-E_{0,\ell})}{E_{b}}}\right)}{\cosh\!\left(2\pi\,\sqrt{\frac{2E_{i}}{E_{b}}}\right)-\cosh\!\left(2\pi\,\sqrt{\frac{2(E_{i}-\omega)}{E_{b}}}\right)}$
$\approx e^{-\beta E_{0,\ell}}\,\left(\frac{e^{\beta\omega}}{e^{\beta\omega}-1}-\frac{e^{-\beta\omega}}{1-e^{-\beta\omega}}\right)=e^{-\beta E_{0,\ell}}\approx 1\,.$ (3.23)
Then $\mathcal{A}^{\Delta}_{E_{i}+\omega}-\mathcal{A}^{\Delta}_{E_{i}-\omega}$ is given by (3.20) and the semiclassical greybody factor reduces to
$P_{\text{gb}}\approx|\mathcal{N}_{\ell}|^{2}\,\,\frac{2\,(2\pi)^{2\Delta}\,(2\ell+1)}{\Gamma(2\Delta)}\,\frac{r_{0}^{2(\Delta+\ell)}\,\omega^{2\ell+1}}{\beta^{2\Delta-1}}\,\sinh\!\left(\!\frac{\beta\omega}{2}\!\right)\,\Gamma\!\left(\!\Delta\pm i\,\frac{\beta\omega}{2\pi}\!\right)\,.$ (3.24)
@@
It is also illustrative to take the very low-frequency limit $\omega\ll\beta^{-1}$. Our result (3.24) simplifies to
@@ 4.1 Quantum spin transparency
In this case there is no obstacle to absorption for any frequency $\omega>0$, while stimulated emission will occur only if it takes the black hole down to a state above the spin gap, i.e., $\omega<E_{i}-E_{0,\ell}$. When no spin is involved, i.e., $\ell=0$, stimulated emission is restricted to $\omega<E_{i}$, which leads to the kinks in Figure 1 of Emparan:2025sao for the absorption cross-section as a function of $\omega$. [^10] We take this opportunity to correct a typo in Figs. 2 and 3 of Emparan:2025sao : the caption states $E_{i}=0.1\,E_{b}$, but the plots actually correspond to $E_{i}=0.01\,E_{b}$.
@@
In the analysis and figures below, we focus on this regime with $\ell\geq 1$ and $E_{i}<E_{0,\ell}$, where the distinctive quantum features of transparency and absence of stimulated emission emerge.
The spin-induced gap also closes the channel for single-quantum emission, suppressing spontaneous emission. This led Brown:2024ajk to consider spin-singlet pairs of entangled photons or gravitons, whose spontaneous emission typically produces a slight difference between ingoing and outgoing fluxes. In our setup, however, the fixed-helicity radiation state has zero overlap with such di-photon or di-graviton states, so there is no absorption or stimulated emission of these di-quanta within the window (4.5), and the absorption cross-section vanishes exactly.
Interestingly, a generic coherent wave—such as an unpolarized electromagnetic or gravitational wave with support across multiple angular momenta $\ell$—will generally have nonzero overlap with di-quanta states. The same applies to linearly polarized waves. In such cases, absorption and stimulated emission of di-quanta occur, and the black hole is no longer perfectly transparent. These second-order processes are, however, strongly suppressed Brown:2024ajk , so the black hole remains much fainter in the window (4.5) than in the regime (4.1) above the spin-induced gap.
@@ 4.2 Results for $\ell=1$ photon and $\ell=2$ graviton
$\Bigg[$ $\frac{\left(4E_{b}^{2}+4E_{b}(2E_{i}+\omega)+\omega^{2}\right)\left(E_{b}^{2}+4E_{b}(2E_{i}+\omega)+4\omega^{2}\right)\sinh\!\left(\!2\pi\sqrt{\frac{2(E_{i}+\omega-E_{b})}{E_{b}}}\right)\Theta(E_{i}+\omega-E_{b})}{\cosh\!\left(\!2\pi\sqrt{\frac{2(E_{i}+\omega)}{E_{b}}}\right)-\cosh\!\left(\!2\pi\sqrt{\frac{2E_{i}}{E_{b}}}\right)}$
$-\frac{\left(4E_{b}^{2}+4E_{b}(2E_{i}-\omega)+\omega^{2}\right)\left(E_{b}^{2}+4E_{b}(2E_{i}-\omega)+4\omega^{2}\right)\sinh\!\left(\!2\pi\sqrt{\frac{2(E_{i}-\omega-E_{b})}{E_{b}}}\right)\Theta(E_{i}-\omega-E_{b})}{\cosh\!\left(\!2\pi\sqrt{\frac{2E_{i}}{E_{b}}}\right)-\cosh\!\left(\!2\pi\sqrt{\frac{2(E_{i}-\omega)}{E_{b}}}\right)}\Bigg]$ (4.6)
@@
$\Bigg[$ $\frac{\left(E_{b}^{2}+4E_{b}(2E_{i}+\omega)+4\omega^{2}\right)\sinh\!\left(\!2\pi\sqrt{\frac{2(E_{i}+\omega-3E_{b})}{E_{b}}}\right)\Theta(E_{i}+\omega-3E_{b})}{\cosh\!\left(\!2\pi\sqrt{\frac{2(E_{i}+\omega)}{E_{b}}}\right)-\cosh\!\left(\!2\pi\sqrt{\frac{2E_{i}}{E_{b}}}\right)}$
$-\frac{\left(E_{b}^{2}+4E_{b}(2E_{i}-\omega)+4\omega^{2}\right)\sinh\!\left(\!2\pi\sqrt{\frac{2(E_{i}-\omega-3E_{b})}{E_{b}}}\right)\Theta(E_{i}-\omega-3E_{b})}{\cosh\!\left(\!2\pi\sqrt{\frac{2E_{i}}{E_{b}}}\right)-\cosh\!\left(\!2\pi\sqrt{\frac{2(E_{i}-\omega)}{E_{b}}}\right)}\Bigg]\,.$ (4.7)
@@
Figure 3: Field-type-independent absorption cross-section $\widehat{\sigma}$ (cf. (4.8)) as a function of $\omega$ for $E_{i}=0.8\,E_{b}$. Left: plot at fixed $\ell=1$ for different $\Delta$. Right: plot at fixed $\Delta=3$ for different $\ell$. In the quantum regime where $\omega/E_{i}\lesssim\mathcal{O}(1)$ (say, $\omega/E_{i}\leq 5$), but above the transparency window for the $\ell=2$ graviton, which ends at $\omega/E_{i}=2.75$, the cross-section becomes higher as we increase $\Delta$, and as we decrease $\ell$. The same qualitative feature is observed for fixed $\ell=2$, or for fixed $\Delta=2$. The length unit for $\widehat{\sigma}$ is $E_{b}^{\Delta+\ell-1}\,r_{0}^{\Delta+\ell}$.